ABC Inc. is a mature manufacturing firm. The company just paid dividends of $7.88 per share, but management expects that dividends will decrease by 3.8 percent per year indefinitely. If you require a rate of return of 11.7 percent on this stock, what will you pay for a share today? A. $53 B. $45 C. $41 D. $49
Answer #1

Current price=D1/(Required return-Growth rate)

=(7.88*(1-0.038))/(0.117-(0.038)

=7.58056/0.155

=$49(Approx).